1. The Pizza Place delivers to any location within a radius of 5 miles from the store for free. A delivery person drives 32 blocks north and then 45 blocks east to deliver a pizza. In this city, there are about 6 blocks per half mile.
  2. a. Should there be a charge for delivery? Explain. b. Describe two delivery situations that would
  3. result in about 5 miles. 
1
Expert's answer
2021-08-06T10:25:40-0400

"\text{From the question, we have that the resultant distance is }\\\sqrt{45^2+32^2}=\text{55.22 blocks}\\\text{Since, there are six blocks per half mile, automatically there are 12 blocks per mile,}\\\text{therefore there are $\frac{55.22}{12}=4.6miles$}
\\\text{Therefore there should be no charge for the delivery since it is within the 5 miles range}\\\text{If the delivery point is 40 blocks north and 50 blocks west, we have that the resultant distance}\\\text{is 64.03 blocks which is 5.34miles} 
\\\text{If he also moves in the opposite direction that is 40 blocks north and 50 blocks east, the }\\\text{resultant distance exceeds 5 miles}"